Let VPS =8V and VT=0.7V ,Assuming rf=0. calculate the value of R such that ID=1.20mA, if Vps is reduced to 4v and ohms is charged to 3.5k. determine the power dissipated in the diode
1
Expert's answer
2020-07-07T07:35:54-0400

Vps= 8 volt , VT= 0.7 volt Rf= 0,Id=1.20 mA

R= (8-0.7)/1.2= 6.08 K ohm

Vps=4 volt ,R= 3.5 Kohm

P= 16/3.5=4.57 mwatt